# New Foundations is consistent

### v22

@@ 3.3 Hypotheses for the recursion
    We further intimate that for each $x\in\tau_{\gamma}$, $-1\leq\gamma<\alpha$, we have defined objects $S$ for which we say that $S$ is a support of $x$. [This is introduced before the definition of support is actually given].
@@
    We further stipulate that extensionality holds for each $\beta\in\alpha$ (for each $\gamma\in\beta$, any $x\in\tau_{\beta}$ is uniquely determined by $x\cap\tau_{\gamma}$; $x$ is uniquely determined by $x\cap\tau_{-1}$ only on the additional assumption that $x\cap\tau_{-1}$ is nonempty).
@@
    We assert that $\iota_{*}(N)\leq\iota_{*}^{+}(N_{\delta},S)$ will hold for any near litter $N$ and support $S$ of $N_{\delta}$, $\delta<\alpha$. This depends on the definition of the position function $\iota_{*}$ which appears below at a convenient point but involves no recursion. (I7) appears on a list of hypotheses about $\iota_{*}^{+}$ which appears later, because it is used early.
@@ 3.4 Machinery for enforcing extensionality in the model
We further stipulate (I6) that extensionality holds for each $\beta\in\alpha$ (for each $\gamma\in\beta$, any $x\in\tau_{\beta}$ is uniquely determined by $x\cap\tau_{\gamma}$; $x$ is uniquely determined by $x\cap\tau_{-1}$ only on the additional assumption that $x\cap\tau_{-1}$ is nonempty).
@@
The function may be constructed directly: first choose an ordering of type $\mu$ on litters, then put singletons of atoms directly after the litter they are inside (and before all later litters), then put near-litters after the corresponding litter and all singletons of atoms in the symmetric difference, then map each near-litter or singleton of an atom to its position in the resulting order. We don’t run out of room before finishing the construction because $\mu$ has cofinality at least $\kappa$.
We give a detailed description of how to do this. Choose a well-ordering $\leq_{0}$ of order type $\mu$ on the set of all near-litters.
We refer to a near-litter which is not a litter as a proper near-litter. With each litter, singleton of atom, or proper near-litter associate a triple of ordinals. For a litter, the first ordinal will be its position in the order $\leq_{0}$, and the next two will be 0. For a singleton of an atom, the first ordinal will be the position in $\leq_{0}$ of the litter $L$ which includes it, the second will be the successor of the position of its element in $\leq_{L}$, and the third will be zero. For a proper near-litter $N$, the first ordinal will be the supremum of the position in $\leq_{0}$ of $N^{\circ}$ and the positions in $\leq_{0}$ of the litters in containing each element of $N\Delta N^{\circ}$ ($<\mu$ by cofinality), the second ordinal will be the supremum of the positions of the elements $x$ of $N\Delta N^{\circ}$ in the respective orders $\leq_{x^{\circ}}$, where $x^{\circ}$ is the litter containing $x$ ($<\kappa$ by regularity), and the third will be the successor of the position of $N$ itself in $\leq_{0}$.
The order on the litters, singletons of atoms, and proper near-litters will be determined by lexicographic order on the associated triples. This is clearly a well-ordering designed explicitly to have the properties indicated above, and has order type $\mu$ because each item in this order can be seen to have $<\mu$ possible predecessors, on the basis of its first coordinate: this is obvious for a triple representing a litter; for a triple representing a singleton of an atom, there are $\kappa$ choices for the second coordinate and one for the third; for a triple representing a near-litter, there are $\kappa$ choices for the second coordinate, and there are $<\mu$ near litters which are subsets of the union of the litters with position in $\leq_{0}$ at or before the first coordinate, simply because this union is a set of size $<\mu$ which has $<\mu$ subsets because $\mu$ is strong limit, so there are $<\mu$ possible values for the third coordinate.
@@
So, what remains is the case of $x$ with $x\cap\tau_{-1}$ empty and each $x\cap\tau_{\gamma}$ nonempty for $\gamma<\beta$.
If $x\cap\tau_{\gamma}$ is an extension of $x$ and $a\in x\cap\tau_{\gamma}$ is not of the form $N_{\gamma}$ for a near-litter $N$, then $x\cap\tau_{\gamma}$ must be the unique distinguished extension: the reason for this is that for any distinguished extension, all elements $b\in x\cap\tau_{\delta}$ of other extensions must be $N_{\delta}$’s.
So we are down to the case where $x\cap\tau_{\delta}$ is nonempty for $\delta\neq-1$, $x\cap\tau_{-1}$ is empty, and each element of any $x\cap\tau_{\delta}$ is of the form $N_{\delta}$ where $N$ is a near-litter. Let $x\cap\tau_{\gamma}$ be a distinguished extension. For any $M_{\delta}\in x$ with $\delta\neq\gamma$, we have $M\sim f_{\gamma,\delta}(P_{\gamma},Q)$ for some near-litter $P$ and support $Q$ of $P_{\gamma}$. Then we have $\iota_{*}(M)>\iota^{+}_{*}(P_{\gamma},Q)\geq\iota_{*}(P)$ by (I7). Let $N_{\gamma}\in x$ be chosen so that $\iota_{*}(N)$ is minimal. It follows that for any $\delta\neq\gamma$ and $M_{\delta}\in x$ we have $\iota_{*}(M)>\iota_{*}(N)$, from which it is evident that we cannot have two distinct distinguished extensions: if $\delta$ were the index of another distinguished extension, and $\iota_{*}(M)$ were chosen minimal so that $M_{\delta}\in x$, it would follow that $\iota_{*}(M)<\iota_{*}(N)$ as above but also that $\iota_{*}(N)<\iota_{*}(M)$, which is absurd. ∎
Definition 3.14 ($A$ map).
Note that the distinguished extension of any type element $x$ is the image under $A^{-1}$ of its other extensions.
Proposition 3.15.
No subset of a type has infinitely many iterated images under $A^{-1}$.
Proof.
Let $a$ be a subset of type $\gamma$ for which $A^{-1}(a)$ exists.
Since $A^{-1}(a)$ exists, every element of $a$ is of the form $N_{\gamma}$ where $N$ is a near-litter. Choose $N_{\gamma}\in A^{-1}(a)$ such that $\iota_{*}(N)$ is minimal. Note that $N$ is in fact a litter by the constraints in definition 3.10.
Let $b=A^{-1}(a)$: we have $a=A_{\gamma}(b)$, where $b$ is a subset of some $\tau_{\delta}$ with $\delta\neq\gamma$. In particular, $N=f_{\delta,\gamma}(u,U)$ for some $u\in\tau_{\delta}$ and $U$ a support of $u$. If $u=M_{\delta}$ we can further state that $\iota_{*}(N)>\iota^{+}_{*}(M_{\delta},U)\geq\iota_{*}(M)$: if $b$ itself has an image under $A^{-1}$, the minimum value of ordinals $\iota_{*}(M)$ for $M_{\delta}\in b$ will be less than the minimum value of ordinals $\iota_{*}(N)$ for $N_{\gamma}\in a$, which establishes that there is an ordinal parameter determined by a nonempty subset of a type which decreases strictly when $A^{-1}$ is applied (if it is applicable), and so no nonempty subset of a type can have infinitely many iterated preimages under $A^{-1}$. This is a rephrasing of an argument which occurred above in the discussion of the uniqueness of distinguished extensions. ∎
@@
We say that an element of a type is extensional iff it is pre-extensional and its distinguished extension has an even number of iterated images under $A^{-1}$. This implies that each of its other extensions has an odd number of iterated images under $A^{-1}$. [^4] We do know that we are carefully, explicitly, spelling out a construction which looks very much like the construction of the bijection in the Cantor-Schröder-Bernstein theorem. But the details of the maps involved are used, so everything must be spelled out.
@@ 3.5 Allowable permutations and supports
More generally, for any finite subset $A$ of $\lambda\cup\{-1\}$ with maximum $\alpha$, define $\pi_{A}$ as $(\pi_{A\setminus\{{\tt min}(A)\}})_{{\tt min}(A)}$ and $\pi_{\{\alpha\}}=\pi$. The maps $\pi_{A}$ may be referred to as iterated derivatives of $\pi$ [^5] There is a silly notational point here: we might want to suppose $\pi_{A}$ and $\pi_{\alpha}$ to be essentially distinguished in some way we do not actually implement (for example by type face) in order to prevent confusion of $\pi_{\{0\}}$ with $\pi_{1}$. A similar problem exists due to the identification of finite ordinals with finite sets of ordinals. However, it can also be noted that $\pi_{\{0,\ldots,n\}}$ and $\pi_{n+1}$ cannot make sense for the same allowable permutation $\pi$, so we think this is harmless. It should be clear that a structural permutation is exactly determined by its iterated derivatives which are $-1$-structural.
@@ A $\beta$-support condition ($-1\leq\beta\leq\alpha$) is defined as a pair $(x,A)$, where
We may write $S_{\delta}$ intead of $S(\delta)$.
@@
    We define the action of a $\beta$-allowable permutation $\pi$ on a $\beta$-support $S$: if $S(\delta)=(x,A)$, $\pi[S](\delta)=(\pi_{A}``x,A)$.
@@
    An element $x$ of $\tau^{*}_{\beta}$ has $\beta$-support $S$ iff for every $\beta$-allowable permutation $\pi$, if $\pi[S]=S$ then $\pi(x)=x$. We say that an element $x$ of $\tau^{*}_{\beta}$ which has a $\beta$-support is $\beta$-symmetric.
@@
The motivation for this is that we need $\beta$-allowable permutations ($\beta\leq\alpha$) to send extensional elements of supertypes to extensional elements. Suppose $x\in\tau_{\beta}$ and $x\cap\tau_{\gamma}=\{b\}$. If $x$ is extensional, this has to be the distinguished extension of $x$. For any $\delta\in\beta\setminus\{\gamma\}$, it follows that $x\cap\tau_{\delta}$ is the set of all $N_{\delta}$ such that $N\sim f_{\gamma,\delta}(b,S)$ for some support $S$ of $b$. This tells us that a $\beta$-allowable permutation $\pi$, such that $\pi(x)$ has $\gamma$-extension $\{\pi_{\gamma}(b)\}$, must have the $\delta$-extension of $\pi(x)$ equal to
@@
for each support $S$ of $b$. The coherence condition enforces this neatly, showing that it is motivated by considerations required to get extensionality to work: the action of $\pi_{\beta}$ conveniently correlates supports of $b$ with supports of $\pi_{\beta}(b)$.
@@
Recall that we defined $A_{\delta}(a)$ in definition 3.14 as
@@ 3.6 Model elements defined
Remark 3.28 ($\kappa$-completeness of the structure).
For any subset $X$ with cardinality $<\kappa$ of a type $\gamma$ and $\beta>\gamma$, it should be clear that $X_{\beta}$ has a support, whose range is obtained from the union of the ranges of the supports of the elements of $X$ by replacing each element $(u,B)$ of the union of the ranges with $(u,B\cup\{\beta\})$, and therefore belongs to the model. $X_{\beta}$ is obviously extensional (the extension $X$ is clearly the distinguished extension and has no image under $A^{-1}$).
Definition 3.29 (position functions).
The conditions constraining the choice of functions $\iota^{+}_{\beta}$ (enforcing (I4), (I7), (I11)) are
@@
There is no difficulty in satisfying these constraints given that $\tau_{\beta}^{+}$ has size $\mu$ (I2) as there are only a small set of constraints on any particular value of $\iota_{*}^{+}$ and $\mu$ is of cofinality at least $\kappa$.

### v23

@@ 3.3 Hypotheses for the recursion
This set of inductive hypotheses is not minimal, but the way it is presented has been chosen to match its actual usage in the following sections.
@@
    We further intimate that for each $x\in\tau_{\gamma}$, $-1\leq\gamma<\alpha$, we have defined objects $S$ for which we say that $S$ is a $\gamma$-support of $x$. [This is introduced before the definition of support is actually given in 3.22, but there is no technical obstacle to defining supports earlier if desired].
@@
    We further stipulate that extensionality holds for each $\beta\in\alpha$: that is, for each $\gamma\in\beta$, any $x\in\tau_{\beta}$ is uniquely determined by $x\cap\tau_{\gamma}$; $x$ is uniquely determined by $x\cap\tau_{-1}$ only on the additional assumption that $x\cap\tau_{-1}$ is nonempty.
@@
    We assert that $\iota_{*}(N)\leq\iota_{*}^{+}(N_{\delta},S)$ will hold for any near litter $N$ and support $S$ of $N_{\delta}$, $\delta<\alpha$. This depends on the definition of the position function $\iota_{*}$ which appears below at a convenient point but involves no recursion. (I7) appears in (I11), but is mentioned earlier here because it is used earlier in the argument.
@@ 3.4 Machinery for enforcing extensionality in the model
We further stipulate (I6) that extensionality holds for each $\beta\in\alpha$: that is, for each $\gamma\in\beta$, any $x\in\tau_{\beta}$ is uniquely determined by $x\cap\tau_{\gamma}$; $x$ is uniquely determined by $x\cap\tau_{-1}$ only on the additional assumption that $x\cap\tau_{-1}$ is nonempty.
@@
The function may be constructed directly: first choose an ordering of type $\mu$ on litters, then put singletons of atoms directly after the litter they are inside (and before all later litters), then put near-litters after the corresponding litter and all singletons of atoms in the symmetric difference, then map each near-litter or singleton of an atom to its position in the resulting order. We do not run out of room before finishing the construction because $\mu$ has cofinality at least $\kappa$. We prove more formally that there are such functions just below.
Lemma 3.11 (position function generator).
If $X$ is a set of size at most $\mu$ and $f:X\to\mathcal{P}(\mu)$ is such that each $f(x)$ is bounded, then there is an injection $\iota:X\to\mu$ such that $\iota(x)\not\in f(x)$.
Proof.
Choose a well-ordering $<_{X}$ of $X$ of type at most $\mu$. At each stage $x$, assume we have already constructed $\iota(y)$ for $y<_{X}x$. We then define
$\iota(x)=\min\left(\mu\setminus f(x)\setminus\{\iota(y):y<x\}\right).$
Note that as $f(x)$ is bounded in $\mu$, and $\mu$ has initial order type, the set $\mu\setminus f(x)$ is of cardinality $\mu$. Since $\{\iota(y):y<x\}$ has cardinality less than $\mu$, the difference must be nonempty, so the definition given always succeeds.
We now show how to construct a position function for near-litters and singletons of atoms satisfying the given constraints.
Proof.
First, let $\iota_{0}$ be any injection from the set of litters into $\mu$. Next, use lemma 3.11, with $X$ taken to be the collection of all singletons of atoms, setting $f(\{a\})=\{\nu:\nu\leq\iota_{0}(L)\}$ where $L$ is the unique litter such that $a\in L$, to create $\iota_{1}$.
Next, use lemma 3.11 again, with $X$ taken to be the collection of all near-litters, setting
$f(N)=\{\nu:\nu\leq\iota_{0}(N^{\circ})\}\cup\bigcup_{a\in N\Delta N^{\circ}}\{\nu:\nu\leq\iota_{1}(\{a\})\},$
to obtain $\iota_{2}$.
Finally, we define a map $\iota$ from near-litters and singletons of atoms to $3\times\mu$ by
$g(x)=\begin{cases}(0,\iota_{0}(x))&\text{if $x$ is a litter}\\ (1,\iota_{1}(x))&\text{if $x$ is a singleton}\\ (2,\iota_{2}(x))&\text{otherwise}\end{cases}$
We can then define $\iota_{*}(x)$ to be the position of $g(x)$ in the restriction of the lexicographic order on $3\times\mu$ to the image of $g$ (which is of order type $\mu$).
The resulting function $\iota_{*}$ is then clearly a bijection, and has the required properties by construction, noting that near-litters that happen to be litters will automatically satisfy the third condition.
@@
We are now going to describe how extensionality is enforced in our model. Each model element will have a particular extension, called its distinguished extension, from which all alternate extensions can be calculated using the $f$ maps.
@@
Definition 3.14 (alternate extension map).
Note that any distinguished extension of any type element $x$ is the image under $A^{-1}$ of its other extensions.
Using properties of the $f$ maps, we can prove the following useful well-foundedness constraint.
Proposition 3.15.
No nonempty subset of a type has infinitely many iterated images under $A^{-1}$, and there are no cycles in $A^{-1}$.
Proof.
Let $a$ be a nonempty subset of type $\gamma$ for which $A^{-1}(a)$ exists.
Since $A^{-1}(a)$ exists, every element of $a$ is of the form $N_{\gamma}$ where $N$ is a near-litter. Choose $N_{\gamma}\in a$ such that $\iota_{*}(N)$ is minimal. Note that $N$ is in fact a litter by the constraints in definition 3.10.
Let $b=A^{-1}(a)$: we have $a=A_{\gamma}(b)$, where $b$ is a subset of some $\tau_{\delta}$ with $\delta\neq\gamma$. In particular, $N=f_{\delta,\gamma}(u,U)$ for some $u\in\tau_{\delta}$ and $U$ a support of $u$. If $u=M_{\delta}$ we can further state that $\iota_{*}(N)>\iota^{+}_{*}(M_{\delta},U)\geq\iota_{*}(M)$: if $b$ itself has an image under $A^{-1}$, the minimum value of ordinals $\iota_{*}(M)$ for $M_{\delta}\in b$ will be less than the minimum value of ordinals $\iota_{*}(N)$ for $N_{\gamma}\in a$, which establishes that there is an ordinal parameter determined by a nonempty subset of a type which decreases strictly when $A^{-1}$ is applied (if it is applicable), and so no nonempty subset of a type can have infinitely many iterated images under $A^{-1}$, nor can there be any cycles in $A^{-1}$. ∎
@@
So, what remains is the case of $x$ with $x\cap\tau_{-1}$ empty and each $x\cap\tau_{\gamma}$ nonempty for $\gamma<\beta$. Suppose that $x\cap\tau_{\gamma}$ and $x\cap\tau_{\delta}$ were both distinguished. Then we would have
$x\cap\tau_{\gamma}=A^{-1}(x\cap\tau_{\delta})=A^{-1}(A^{-1}(x\cap\tau_{\gamma})),$
which contradicts proposition 3.15 as this would imply that $A^{-1}$ has a cycle of length 2. ∎
@@
We say that an element of a type is extensional iff it is pre-extensional and its distinguished extension has an even number of iterated images under $A^{-1}$. This implies that each of its other extensions has an odd number of iterated images under $A^{-1}$. [^8] We do know that we are carefully, explicitly, spelling out a construction which looks very much like the construction of the bijection in the Cantor–Schröder–Bernstein theorem. But the details of the maps involved are used, so everything must be spelled out.
@@ 3.5 Allowable permutations and supports
More generally, for any finite subset $A$ of $\lambda\cup\{-1\}$ with maximum $\alpha$ and for any $\alpha$-structural permutation $\pi$, define $\pi_{A}$ as $(\pi_{A\setminus\{{\tt min}(A)\}})_{{\tt min}(A)}$ and $\pi_{\{\alpha\}}=\pi$. The maps $\pi_{A}$ may be referred to as iterated derivatives of $\pi$ [^9] There is a silly notational point here: we might want to suppose $\pi_{A}$ and $\pi_{\alpha}$ to be essentially distinguished in some way we do not actually implement (for example by type face) in order to prevent confusion of $\pi_{\{0\}}$ with $\pi_{1}$. A similar problem exists due to the identification of finite ordinals with finite sets of ordinals. However, it can also be noted that $\pi_{\{0,\ldots,n\}}$ and $\pi_{n+1}$ cannot make sense for the same allowable permutation $\pi$, so we think this is harmless. It should be clear that a structural permutation is exactly determined by its iterated derivatives which are $-1$-structural.
@@
We will now define a support to be a particular kind of object that can be used to track how an allowable permutation acts along different derivatives on a small part of our model.
@@ A $\beta$-support condition ($-1\leq\beta\leq\alpha$) is defined as a pair $(x,A)$, where
We may write $S_{\delta}$ instead of $S(\delta)$.
@@
    We define the action of a $\beta$-allowable permutation $\pi$ on a $\beta$-support $S$: if $S(\delta)=(x,A)$, $\pi[S](\delta)=(\pi_{A}``x,A)$. This means that $\pi[S]=S$ iff $\pi_{A}``x=x$ for all $(x,A)\in{\tt rng}(S)$.
@@
    An element $x$ of $\tau^{*}_{\beta}$ has $\beta$-support $S$ iff for every $\beta$-allowable permutation $\pi$, if $\pi[S]=S$ then $\pi(x)=x$. We say that an element $x$ of $\tau^{*}_{\beta}$ which has a $\beta$-support is $\beta$-symmetric. We allow ourselves to say briefly “$x$ has support $S$” and “$x$ is symmetric” if $x\in\tau_{\beta}$ and $S$ is a $\beta$-support of $x$.
@@
Recall the coherence condition for $\beta$-allowable permutations ($\beta\leq\alpha$): for any $(x,S)\in\tau_{\gamma}^{+}$ and any $\delta\in\beta\setminus\{\gamma\}$,
The motivation for this is that we need $\beta$-allowable permutations to send extensional elements of supertypes to extensional elements. Suppose $x\in\tau_{\beta}$ and $x\cap\tau_{\gamma}=\{b\}$. If $x$ is extensional, this has to be the distinguished extension of $x$. For any $\delta\in\beta\setminus\{\gamma\}$, it follows that $x\cap\tau_{\delta}$ is the set of all $N_{\delta}$ such that $N\sim f_{\gamma,\delta}(b,S)$ for some support $S$ of $b$. This tells us that a $\beta$-allowable permutation $\pi$, such that $\pi(x)$ has $\gamma$-extension $\{\pi_{\gamma}(b)\}$, must have the $\delta$-extension of $\pi(x)$ equal to
@@
for each support $S$ of $b$. The coherence condition enforces this neatly, showing that it is motivated by considerations required to get extensionality to work: the action of $\pi_{\gamma}$ conveniently correlates supports of $b$ with supports of $\pi_{\gamma}(b)$.
@@
Recall that for nonempty $a\subseteq\tau_{\gamma}$ we defined $A_{\delta}(a)$ in definition 3.14 as
@@ 3.6 Model elements defined
Proposition 3.29 (symmetric comprehension).
Let $X$ be a subset of a type $-1<\beta<\alpha$. Suppose that there is an $\alpha$-support $S$ with the property that for any $\alpha$-allowable $\pi$, when $\pi[S]=S$, we have $\pi_{\beta}``X=X$. Then $X=x\cap\tau_{\beta}$ for some (necessarily unique) model element $x\in\tau_{\alpha}$.
Proof.
If $X$ is empty, then $X$ is simply the $\beta$-extension of the empty model element. Suppose $X$ is nonempty, and let $n$ be the number of iterated images that $X$ has under $A^{-1}$.
If $n$ is even, consider the element $x\in\tau^{*}_{\alpha}$ such that $x\cap\tau^{*}_{\beta}=X$, $x\cap\tau^{*}_{-1}=\emptyset$, and whose other extensions at levels $-1<\gamma<\alpha$, $\gamma\neq\beta$ are given by $x\cap\tau^{*}_{\gamma}=A_{\gamma}(X)$. We show this is a model element. Clearly each extension satisfies $x\cap\tau^{*}_{\gamma}\subseteq\tau_{\gamma}$. It is pre-extensional because $x\cap\tau_{\beta}$ is its distinguished extension, and it is extensional as $n$ is even. It has support $S$: if $\pi$ is $\alpha$-allowable and fixes $S$, then $\pi_{\beta}``(x\cap\tau_{\beta})=x\cap\tau_{\beta}$ by assumption, so by extensionality, $\pi$ must fix $x$. So $x\in\tau_{\alpha}$ and $x\cap\tau_{\beta}=X$.
If $n$ is odd, let $Y=A^{-1}(X)\subseteq\tau_{\gamma}$. We claim that $S$ has the property that if $\pi$ is $\alpha$-allowable and $\pi[S]=S$, then $\pi_{\gamma}``Y=Y$. Indeed, as allowable permutations commute with $A$ maps by proposition 3.27,
$\pi_{\gamma}``Y=\pi_{\gamma}``A^{-1}(X)=A^{-1}(\pi_{\beta}``X)=A^{-1}(X)=Y.$
This allows us to repeat the argument of the previous paragraph using $Y$ to obtain a model element $x\in\tau_{\alpha}$ satisfying $x\cap\tau_{\gamma}=Y$, where $x\cap\tau_{\gamma}$ is distinguished. Then, by pre-extensionality, we must have $x\cap\tau_{\beta}=A_{\beta}(x\cap\tau_{\gamma})=X$, as required. ∎
Remark 3.30.
All model elements for $\alpha>0$ are of this form: if $x\in\tau_{\alpha}$ and $\beta\in\alpha$, then any support $S$ for $x$ satisfies the hypothesis that $\pi[S]=S$ implies $\pi_{\beta}``(x\cap\tau_{\beta})=x\cap\tau_{\beta}$.
Corollary 3.31.
The singleton $\{x\}_{\alpha}$ exists for all $x\in\tau_{\beta}$.
Proof.
If $S$ is a support for $x$, then $S^{\uparrow\alpha}$ satisfies the hypothesis of symmetric comprehension as required. ∎
Corollary 3.32 ($\kappa$-completeness of the structure).
Let $\beta\in\alpha$. For any small subset $X$ of type $\alpha$, there is a model element $y\in\tau_{\alpha}$ such that
$y\cap\tau_{\beta}=\bigcup_{x\in X}x\cap\tau_{\beta}.$
Proof.
Let $S$ be a support whose range is obtained from the unions of ranges of supports for each $x\in X$. Then, if $\pi[S]=S$, we have
$\pi_{\beta}``\left(\bigcup_{x\in X}x\cap\tau_{\beta}\right)=\bigcup_{x\in X}\pi_{\beta}``(x\cap\tau_{\beta})=\bigcup_{x\in X}\pi(x)\cap\tau_{\beta}=\bigcup_{x\in X}x\cap\tau_{\beta}.$
The conclusion then follows directly from symmetric comprehension. ∎
Corollary 3.33.
As small unions exist and singletons exist, $X_{\alpha}$ must also exist for any small set $X\subseteq\tau_{\beta}$.
Proposition 3.34 (position functions).
Given that $\tau_{\beta}^{+}$ has size $\mu$ (I2), there exist a choice of injections $\iota^{+}_{\beta}$ enforcing (I4), (I7), (I11). More precisely,
@@
Proof.
We use the position function generator (proposition 3.11): recall that this states that if $X$ is a set of size at most $\mu$ and $f:X\to\mathcal{P}(\mu)$ is such that each $f(x)$ is bounded, then there is an injection $\iota:X\to\mu$ such that $\iota(x)\not\in f(x)$. Setting $X=\tau^{+}_{\beta}$, we define
$g(x,S)=\bigcup_{N\text{ s.t. }x=N_{\delta}}\iota_{*}(N)\cup\bigcup_{(y,A)\in{\tt rng}(S)}\iota_{*}(y)$
and let $f(x,S)=\{\nu:(\exists\xi\in g(x,S):\nu\leq\xi)\}$. As $\mu$ has cofinality at least $\kappa$ and each $g(x,S)$ is small, each $f(x,S)$ is bounded as required. ∎